OpenStudy (rock_mit182):

How many liters of pure water should be mixed with a 8-L solution of 90 % acid to produce a mixture that is 60% water ?

OpenStudy (rock_mit182):

@mathmate @3mar

OpenStudy (rock_mit182):

x= pure water concetration

OpenStudy (rock_mit182):

0.9(8-x )

OpenStudy (3mar):

For the last one, your presentation of the problem is awesome and the formed equation is correct! You are very good! Keep that on! Now for this one!

OpenStudy (rock_mit182):

I don't know if that makes sense..

OpenStudy (rock_mit182):

oh thanks!

OpenStudy (rock_mit182):

yeah now for this one...

OpenStudy (3mar):

90 % acid: means it contains 10 % water or what?

OpenStudy (rock_mit182):

initially i guess, but at the end the C of water must be changed to 60%

OpenStudy (3mar):

I kow, just want to know about the acid

OpenStudy (rock_mit182):

C= X/V

OpenStudy (rock_mit182):

8(0.1) = x right ?

OpenStudy (rock_mit182):

initially 0.8 liters of water

OpenStudy (3mar):

Could you give me just 2 min, please?

OpenStudy (rock_mit182):

0.8 + x = 0.6/x i dont know if that is right..

OpenStudy (rock_mit182):

if I'm right could you medal me, I'll medal you anyway...

OpenStudy (rock_mit182):

is it right ?

OpenStudy (3mar):

d